Write a coding sequence with the variables x=1 y=10 z=0 that equals 123 after you run the program using += or *= or -= commands
Answer #1

First we will make z equal to 3 then x equal to 20 and then y = 100

Then adding all these x,y,z we get 123

Pseudo code:

z+=x // z=z+x so z becomes 0+1 =1

z+=x // z=z+x so z becomes 1+1 =2

z+=x // z=z+x so z becomes 2+1=3

//Now make x =20

x*=y // x=x*y so x becomes 1*10 =10

x+=x // x=x+x so x become 10+10=20

//Now make y as 100

y+=y // y=y*y so z becomes 10*10 =100

//Adding x,y,z

y+=x //y=y+x so y becomes 100+20 =120

y+=z //y=y+z so y = 120+3 = 123

print(y) //gives output 123